# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

### v2

@@ Abstract
  1. Accurate uncertainty quantification of large language models (LLMs) provides credibility measure over their outputs. However, fine-tuned LLMs often struggle with overconfidence in uncertain predictions due to the limitations in the models’ ability to generalize with limited data. Existing parameter efficient fine-tuning (PEFT) uncertainty quantification methods for LLMs focus on post fine-tuning stage and fall short of calibrating epistemic uncertainty. To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates epistemic uncertainty over the space of functions that map input prompts to outputs. We implement UQ4CT during the fine-tuning stage via a mixture-of-experts framework that hierarchically decomposes the functional space. We demonstrate that UQ4CT reduces Expected Calibration Error (ECE) by more than $25\%$ while maintaining high accuracy across $5$ benchmarks. Even under distribution shift, UQ4CT maintains superior ECE performance with high accuracy, showcasing improved generalizability.
11footnotetext: Department of Computer Science and Engineering, University of California San Diego, La Jolla, California, USA22footnotetext: Halıcıoğlu Data Science Institute, University of California San Diego, La Jolla, California, USAfootnotetext: Preprint. Under Review. Correspondence to: Yian Ma <yianma@ucsd.edu>.
@@ 1 Introduction
Quantifying credibility of the outputs has been one of the most important problems around large language models (LLMs) (Chang et al., 2024). In particular, fine-tuned LLMs often struggle with overconfidence in gauging the uncertain of their outputs—oftentimes due to limited training data—and does not reflect the true credibility of the answers that they generate (Xiao et al., 2022a; Zhou et al., 2022; He et al., 2023; Tian et al., 2023; OpenAI, 2023). Such overconfidence has become a crucial problem in safety-related decision making and in fields where data is relatively limited, such as formal proof generation, climate science, and healthcare (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022). Methods that enhance uncertainty quantification of fine-tuned LLMs are sought after for the assurance of trustworthy predictions in applications.
One salient challenge of uncertainty quantification in large models is the trade-off among accuracy, calibration, and tractability. One would first like to maintain the capability of the LLMs and not sacrifice accuracy of the generated outputs. Given this desideratum, a number of existing works focus on perturbing the prompts that are fed into the language models and quantify the uncertainty associated with such perturbations (Hou et al., 2023; Farquhar et al., 2024; Gao et al., 2024). These methods capture the aleatoric uncertainty assuming that the trained model corresponds to the ground-truth data distribution. However, these approaches cannot capture the epistemic uncertainty that arises from the model being trained on limited data. To capture this epistemic uncertainty, one would need to introduce changes or perturbations to the model in accordance with the data distribution. On the tractability side, it is not realistic to consider perturbations to all the model parameters for an LLM (Minaee et al., 2024).
We therefore propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT) to calibrate the epistemic uncertainty in LLMs. Since we only focus on the epistemic uncertainty over the functional space that maps input prompts to the outputs, we no longer need to consider uncertainty in the entire parameter space. Instead, we only need to measure the uncertainty regarding the functional relationships. Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the interesting functional relationships arising from the underlying data.
In particular, we make use of the low rank adaptation (LoRA, Hu et al. (2021a)) ensemble at each layer of the LLM to form the basis functions. We then utilize the mixture of experts (MoE) architecture (Li et al., 2024) to hierarchically combine the basis functions and form the functional space. During fine-tuning, UQ4CT simultaneously learns the parameters in the LoRA experts and the prompt-dependent mixture weights to output the calibrated distribution over the functional space. During inference, the LoRA experts capture different functional relationships in the fine-tuning data throughout training, and the MoE routers dynamically select these functional bases conditioned on the input. The selection process models the functional level epistemic uncertainty, and consequently captures the uncertainty in the output space. We calibrate functional level epistemic uncertainty to align with predictive correctness during training time. This approach significantly improves uncertainty estimations of the model on its predictions without compromising accuracy. To summarize, our contributions include:
@@
  • A novel epistemic uncertainty quantification approach with MoE architecture during fine-tuning to quantify functional level epistemic uncertainty and align with predictive correctness, which mitigates overconfidence issue and improves generalizability.
@@
  • A novel calibration loss function incorporating predictive correctness to calibrate the prompt-dependent LoRA mixture for better uncertainty estimation.
@@
  • More than $25\%$ expected calibration error (ECE) reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task; improved ECE performance under distribution shifts on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks without compromising accuracy.
@@ 2 Preliminaries
2.1 Low-rank Adaptation (LoRA)
@@ 2.2 Mixture of Experts (MoE)
The plain transformer block in a large language model consists of the $q,k,v$ encoding layers ($\text{FFN}_{q,k,v}$), layer norm (LN) and feed forward layer (FFN), together with residual connections. Formally, given $\mathbf{h}_{o}=[s_{0},s_{1},...,s_{n}]$ as the tokenized input text, the output of $\ell$-th layer is generated as:
@@
Let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{1\times d}$ $(1\leq i\leq s,1\leq\ell\leq L)$ denote the output hidden state of the $i$-th token at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension. With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-n gate router $R(\cdot)$ chooses n experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-n}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (3)
@@
$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}R^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}\cdot\mathbf{h}^{\ell}$ (4)
where $\mathbf{W}$ is the pretrained weights of the target fine-tuning layer and $\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}$ is the k-th LoRA expert.
@@
Note that the calculation of $E^{\ell}_{k}$ involves $\mathbf{W}^{\ell}$, which represents the weight of the original ATTN or FFN layer.
2.3 Aleatoric and Epistemic Uncertainties
In machine learning models, uncertainty can be categorized into aleatoric (data-wise) and epistemic (model-wise) uncertainty (Hora, 1996; Hüllermeier and Waegeman, 2021). For LLMs, aleatoric uncertainty arises from the inherently ambiguous and context dependent nature of natural languages where a single phrase or sentence can have multiple valid interpretations in different contexts. Epistemic uncertainty is introduced by the model’s lack of knowledge due to limited learning capabilities, suboptimal modeling or sparse training data.
Epistemic uncertainty is highly related to several well-known limitations of generative models. For example, it has been observed that when an LLM is pretrained on a diverse range of text data, it is generally well-calibrated, i.e. the predicted probability of the next token generally aligns with what is observed in real text. However, after fine-tuning or alignment with human preferences, the calibration error deteriorates (Zhao et al., 2021; Achiam et al., 2023a). A related phenomenon is forgetting, where the performance of a fine-tuned LLM diminishes on tasks outside the scope of the target task (Lin et al., 2023; Luo et al., 2023).
Motivated by these observations, we explore functional-level epistemic uncertainty in generative models and aim to develop metrics that assess model performance on specific problem instances to fine-tune the parameter mixture.
@@ 3 Methodology
The high level goal of UQ4CT is to balance the exploration and exploitation of different LoRA experts during fine-tuning. In particular, we incorporate the functional-level epistemic uncertainty (FEU) to calibrate the prompt-dependent parameter mixture with LoRA MoE.
Assume that our answer $a$ is generated via a mixture of mechanisms or models $M$, conditioning on the input prompt $x$. Assume that $e(a)$ is an embedding of $a$ so that least squares distance is a natural distance on the space of $e(a)$. For an expressive enough model class $\mathcal{M}$, and a calibrated distribution $P(M|x)$ over the model class, we can measure the deviation of the generated answer to the “ideal” one $a_{*}=f_{*}(x)$ as:
$\mathbb{E}_{M\sim P(M|x)}\left[\mathbb{E}_{a\sim P(a|M,x)}\left[\|e(a)-e(a_{*})\|^{2}\right]\right]$
$=\mathbb{E}_{P(M|x)}\underbrace{\left[\mathbb{E}_{P(a|M,x)}\left[\|e(a)-\mathbb{E}_{P(a|M,x)}[e(a)]\|^{2}\right]\right]}_{\text{Aleatoric Uncertainty}}+\underbrace{\mathbb{E}_{P(M|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-e(a_{*})\|^{2}\right]}_{\text{Epistemic Uncertainty}}.$ (6)
Note that we hereby quantify uncertainty as a function of the input prompt $x$, since the distribution of model $M$ conditions on $x$. We hence name the task “functional-level uncertainty quantification”.
3.1 Functional-level Epistemic Uncertainty
Figure 1: Mixture of Experts (MoE) architecture to capture and calibrate functional-level epistemic uncertainty. Experts $B_{1...K}A_{1...K}$ capture functional relationships in the data throughout fine-tuning, the weights $\alpha_{1...K}$ quantify the uncertainty in selecting these functional bases conditioned on the input hidden state $h$, which is the semantic representation of the input token $x$. In the UQ4CT workflow, we align this uncertainty with predictive correctness. When the router makes a correct prediction, the loss reinforces this decision, thereby increasing the router confidence in its selection, which aligns with a lower epistemic uncertainty. Conversely, when the router makes an incorrect prediction, the loss penalizes this selection, potentially causing the router to distribute its probabilities more broadly across experts, which is indicative of higher epistemic uncertainty.
Motivated by the decomposition in Eq. (3) for least squares loss, we consider a general distance $d$, and formally define epistemic uncertainty as the variance caused by model training procedure. Specifically, we focus on the variation introduced in the model fine-tuning stage of LLMs. Mathematically, given prompt $x$, we consider the following definition of epistemic uncertainty:
$\text{Epistemic Uncertainty}=\mathbb{E}_{M\sim P(\cdot|x)}\mathbb{E}_{M^{\prime}\sim P(\cdot|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-\mathbb{E}_{P(a^{\prime}|M^{\prime},x)}[e(a^{\prime})]\|^{2})\right]$ (7)
Here, $a^{\prime}$ and $a$ represents the outputs sampled from the parameter mixtures $M^{\prime}$ and $M$ conditioned on the prompt $x$. The epistemic uncertainty measures the least squares distance between $e(a)$ and $e(a^{\prime})$.
@@ 3.2 Decomposition of the Functional Space
Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the functional relationships in the underlying data. Then the functional space can be decomposed into a mixture of the LoRA ensembles. Naively, this decomposition writes:
$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{1}=1}^{K}\sum_{k^{2}=1}^{K}\cdots\sum_{k^{L}=1}^{K}\alpha_{k^{1},k^{2},\dots,k^{L}}g_{k^{L}}^{L}\left(g_{k^{L-1}}^{L-1}\left(\dots g_{k^{1}}^{1}(\mathbf{x})\right)\right),$ (8)
$g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=E_{k^{\ell}}^{\ell}\left(f_{trans}^{\ell}(\mathbf{h}^{\ell})\right),$ (9)
@@
where $f_{trans}^{\ell}$ represents the necessary non-fine-tuned operations within a transformer block (i.e. layer norm, attention calculation, etc.). This linear decomposition requires examining $K^{L}$ parameters to span the entire functional space.
We instead consider a hierarchical decomposition of $f(\mathbf{x})$:
$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{L}=1}^{K}\alpha^{L}_{k^{L}}g_{k^{L}}^{L}\Bigg{(}\sum_{k^{L-1}=1}^{K}\alpha^{L-1}_{k^{L-1}}g_{k^{L-1}}^{L-1}\Bigg{(}\dots\sum_{k^{1}=1}^{K}\alpha^{1}_{k^{1}}g_{k^{1}}^{1}(\mathbf{x})\Bigg{)}\Bigg{)}.$
Here the output function is determined by the combinations of the weights $\alpha^{\ell}_{k^{\ell}}$ across all the layers. As a result, we only need $K\cdot L$ parameters to represent various compositions of the functional relationships.
@@
To further improve the efficiency of representation, we apply a sparse dynamic routing and choose the mixture weights $\alpha_{k^{\ell}}=R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}$ that depend on the input from the previous layer:
$\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}g^{\ell}_{k^{\ell}}(\mathbf{h}^{\ell}).$ (10)
3.3 Quantifying FEU with LoRA MoE Framework
We quantify the functional-level epistemic uncertainty (FEU) with the MoE architecture, which is represented by the embedding $e(a)$ in Eq. (7). As shown in Figure 1, the LoRA experts that are made of matrices $B^{1...L}_{1...K},A^{1...L}_{1...K}$ in Eq. (1) capture important functional relationships in the data during fine-tuning. We treat these functional relationships represented by the LoRA experts as basis functions. As shown in Eq. (10), conditional on the input prompts, the more complex functional relationships that recursively map inputs to outputs are represented as linear combinations (mixture of experts) of the basis functions, hierarchically through each layer.
Uncertainty quantification in this functional space reduces to quantifying the uncertainty of the weights over the basis functions. As shown in Figure 1, the weights $\alpha_{k}^{\ell}$ from the top-n router dynamically mix these basis functions given the input hidden state. The top-n weights that produce the final output hidden state quantify the functional level epistemic uncertainty in the function selection. With mixture of LoRA experts architecture, we follow routing mechanisms of the MoE layers as in Eq. (3) and (4). Specifically, we employ top-2 gate routers, which chooses $2$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$\vspace{-0.5mm}R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (11)
Given an input prompt $x$ with length $s$, we model functional-level epistemic uncertainty (FEU) by aggregating $R^{\ell}(\mathbf{h}^{\ell}_{i})$ over both layer dimension and sequence dimension:
$\text{FEU}(x)=\frac{1}{s}\sum^{s}_{i=1}\left[\frac{1}{L}\sum^{L}_{\ell=1}R^{\ell}(\mathbf{h}^{\ell}_{i})\right].$ (12)
3.4 Training Calibration Loss

### v3

@@ Abstract
  1. Accurate uncertainty quantification in large language models (LLMs) is essential for providing credible confidence estimates over their outputs. However, fine-tuned LLMs often exhibit overconfidence in uncertain predictions, which stems from their limited ability to generalize with sparse data. Existing parameter efficient fine-tuning (PEFT) uncertainty quantification methods for LLMs focus on post fine-tuning stage, and thus fail to address the core issue: limited specialization of PEFT adapters to accurately capture task-specific input-output relationships. To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates uncertainty over the space of functions that map input prompts to outputs. We implement UQ4CT during the fine-tuning stage via a mixture-of-experts framework that hierarchically decomposes the functional space. Empirically, UQ4CT achieves over $25\%$ reduction in Expected Calibration Error (ECE) while preserving high accuracy across five benchmarks. Even under distribution shift, UQ4CT maintains superior ECE performance with high accuracy, showcasing improved generalizability.
11footnotetext: Department of Computer Science and Engineering, University of California San Diego, La Jolla, California, USA22footnotetext: Halıcıoğlu Data Science Institute, University of California San Diego, La Jolla, California, USAfootnotetext: Preprint. Correspondence to: Yian Ma <yianma@ucsd.edu>.
@@ 1 Introduction
Quantifying the credibility of outputs has been one of the most important problems around large language models (LLMs)(Chang et al., 2024). In particular, fine-tuned LLMs often struggle with overconfidence in their outputs due to limited training data, failing to reflect the true credibility of their answers(Xiao et al., 2022; He et al., 2023; Tian et al., 2023; OpenAI, 2023). Such overconfidence can assert misinformation with high certainty, making it difficult for users to discern truth from falsehood. This has become a crucial problem in safety-critical decision making and scientific domains where data is relatively limited, such as formal proof generation, climate science, and healthcare (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022). Methods that enhance uncertainty quantification for fine-tuned LLMs are therefore essential to ensure trustworthy predictions.
One salient challenge of uncertainty quantification in large language models is the trade-off among accuracy, calibration, and efficiency. Ideally, one seeks to calibrate model uncertainty without degrading accuracy or slowing output generation. Recent approaches often focus on prompt perturbation—modifying the model input and quantifying the resulting prediction variance (Hou et al., 2023; Gao et al., 2024)—or sampling multiple completions to measure prediction disagreement (Farquhar et al., 2024). However, these methods generally assume the model is already well-aligned with the data distribution, and thus struggle to capture uncertainty arising during fine-tuning, especially the generalization gap due to adaptation on limited data. Additionally, since these methods require multiple forward passes per input, they incur significant computational overhead, limiting scalability.
Beyond prompt-level approaches, Bayesian methods and ensemble-based uncertainty quantification have been established for fine-tuned LLMs, often in conjunction with low-rank adaptation (LoRA)  (Hu et al., 2021a). Methods such as Monte Carlo dropout (Gal and Ghahramani, 2016), checkpoint ensembles (Chen et al., 2017), deep ensembles (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023), and Laplace-LoRA (Yang et al., 2024a) apply Bayesian inference or ensembling over LoRA parameters to capture uncertainty arising from model adaptation and limited data. While Bayesian and ensemble methods estimate uncertainty after fine-tuning by analyzing the learned parameter space, they do not address the limitations caused by sparse data during fine-tuning. This post hoc perspective can miss uncertainty that arises when adapting to new tasks with limited data.
Figure 1: Left: The Mixture of Experts (MoE) architecture captures diverse functional relationships by dynamically routing inputs to different expert modules based on the prompt. Right: The standard LoRA approach lacks such functional-space diversity, limiting its capacity to capture variations in the functional relationships during fine-tuning.
To overcome this, we shift focus from parameter-space to functional-space uncertainty quantification. The functional space encompasses the input-output mappings the model can realize, capturing the true variability in its predictions. By calibrating uncertainty at this level during fine-tuning, we ensure the model’s confidence better reflects its actual predictive reliability.
We therefore introduce Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), a method that explicitly calibrates functional-level uncertainty in LLMs during fine-tuning. UQ4CT leverages ensembles of LoRA modules at each layer to construct a rich set of basis functions. We then employ a Mixture-of-Experts (MoE) architecture (Li et al., 2024) to hierarchically combine these basis functions, forming a flexible functional space (see Figure 1). During fine-tuning, UQ4CT jointly learns the LoRA expert parameters and calibrates the prompt-dependent function mixture to align functional-level uncertainty with predictive correctness, enabling the model to output calibrated distributions over the functional space.
During inference, LoRA experts offer diverse functional relationships acquired during fine-tuning, while MoE routers dynamically select the most relevant experts for each input prompt. This selection is guided by functional-level uncertainty calibration performed throughout fine-tuning, which aims to optimize the choice of the correct functional relationship for each prompt. More accurate expert selection enables the model to learn diverse functional relationships. As a result, the model’s uncertainty estimates become better aligned, enhancing calibration without compromising accuracy. To summarize, our contributions include:
@@
  • A novel uncertainty quantification approach for LLMs with MoE architecture during fine-tuning to quantify functional-level uncertainty and align with the probability of predictive correctness, which mitigates overconfidence issue and improves generalizability.
@@
  • A new calibration loss that incorporates predictive correctness probability to dynamically align the prompt-dependent LoRA mixture for better uncertainty estimation.
  • Hierarchical decomposition of functional-level uncertainty into layer-wise mixture weights with guarantee that our calibration loss aligns mixture weights with predictive correctness.
@@
  • More than $25\%$ expected calibration error (ECE) reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task; improved ECE performance without compromising accuracy under distribution shift on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks.
@@ 2 Preliminaries
Low-rank Adaptation (LoRA).
@@
The plain transformer block in a large language model consists of the $q,k,v$ encoding layers ($\text{FFN}_{q,k,v}$), layer norm (LN) and the feedforward layer (FFN), together with residual connections. Formally, given $\mathbf{h}^{1}$ as the tokenized input text, the output of $\ell$-th layer is generated as:
@@
Let $\mathbf{h}^{\ell}\in\mathbb{R}^{1\times d}$ $(1\leq\ell\leq L)$ denote the output hidden state at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension. With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-k gate router $\tilde{R}(\cdot)$ chooses $k$ experts with highest probability given a hidden state $\mathbf{h}^{\ell}$:
$\tilde{R}^{\ell}(\mathbf{h}^{\ell})=\text{Keep-Top-k}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}^{\ell})).$ (3)
@@
$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}\tilde{R}^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}_{pre}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}\cdot\mathbf{h}^{\ell}$ (4)
where $\mathbf{W}^{\ell}_{pre}$ is the frozen pretrained weight at layer $\ell$ and $\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}$ is the k-th LoRA expert.
@@ 3 Methodology
The high-level goal of UQ4CT is to leverage the ensemble of prompt-dependent LoRA mixture-of-experts (MoE) to guide and calibrate the confidence of the model during fine-tuning. By quantifying the variability in how different LoRA experts are combined for each input, UQ4CT enables the model to adaptively select expert mixtures that reflect the true uncertainty in its predictions. Our approach not only encourages the model to exploit confident expert combinations for accurate predictions but also promotes exploration of alternative experts when uncertainty is high, ultimately leading to better-calibrated and more reliable model outputs.
Figure 2: MoE architecture to capture functional-level uncertainty. LoRA experts ($B_{i}$, $A_{i}$) capture diverse functional bases, while the top-$k$ router assigns mixture weights based on the input hidden state. The calibration loss aligns functional-level uncertainty with prediction correctness, encouraging confident expert selection for correct predictions and higher uncertainty for incorrect ones.
@@ 3.1 Decomposition of the Functional Space
Given the immense size of both the pre-training dataset and the model, we posit that the pretrained network contains submodules capable of expressing a wide range of functional relationships present in the data. During fine-tuning, our focus is on the functional space spanned by the model, which can be effectively captured as a mixture of LoRA experts, each representing a distinct basis function. However, naively decomposing the full functional space by considering all possible combinations of these experts quickly leads to a combinatorial explosion.
Naive Decomposition. Denote the input prompt as $\mathbf{x}$ and the functional map from input to output as $f$. A straightforward approach is to express the function as a sum over all possible compositions of $K$ submodules per layer, across $L$ layers:
$f(\mathbf{x})=\sum_{k^{1},\cdots,k^{L}=1}^{K}\alpha_{k^{1},\dots,k^{L}}g_{k^{L}}^{L}\bigg{(}\cdots\bigg{(}g_{k^{l}}^{l}\big{(}\dots g_{k^{1}}^{1}(\mathbf{x})\big{)}\bigg{)}\bigg{)},$ (6)
where each $g_{k^{\ell}}^{\ell}$ denotes a particular variant (e.g., LoRA-adapted) of the $\ell$-th block and
$g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=\mathbf{h}^{\ell}+E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\mathbf{h}^{\ell})\big{)},$ (7)
@@
here $f_{trans}^{\ell}$ represents the necessary non-fine-tuned operations within a transformer block (i.e. layer norm, attention calculation, etc.) and as defined in Eq. 4, $E_{k^{\ell}}^{\ell}$ denotes the parameterized adaptation associated with the $k^{\ell}$-th expert in the $\ell$-th layer.
However, this naive decomposition is intractable in practice, as it requires keeping track of $K^{L}$ mixture weights $\alpha_{k^{1},\ldots,k^{L}}$. This is an exponential growth in the number of parameters with respect to both the number of layers $L$ and the number of submodules per layer $K$, which makes the direct approach computationally infeasible for realistic network sizes.
Hierarchical Decomposition. To address the combinatorial explosion of mixture weights, we instead propose a hierarchical decomposition. Here, the mixture at each layer is formed independently, and the output of each layer is a weighted sum over its submodules, with the weights themselves being layer-specific:
$f(\mathbf{x})=\sum_{k^{L}=1}^{K}\alpha^{L}_{k^{L}}g_{k^{L}}^{L}\bigg{(}\cdots\Bigg{(}\sum_{k^{l}=1}^{K}\alpha^{l}_{k^{l}}g_{k^{l}}^{l}\bigg{(}\dots\sum_{k^{1}=1}^{K}\alpha^{1}_{k^{1}}g_{k^{1}}^{1}(\mathbf{x})\bigg{)}\Bigg{)}\Bigg{)}.$ (8)
Instead of needing $K^{L}$ mixture weights, this hierarchical structure only requires $K\cdot L$ weights $\alpha^{\ell}_{k^{\ell}}$, one for each submodule in each layer. This dramatically reduces the parameters required and makes the decomposition tractable, while still enabling a rich set of compositional functions.
@@
Dynamic, Input-Dependent Routing. To further enhance expressivity and efficiency, we allow the mixture weights to depend dynamically on the input at each layer. Specifically, we set the mixture weights to be a sparse routing function $R^{\ell}$ of the hidden state $\mathbf{h}^{\ell}$ at each layer, where $\alpha_{k^{\ell}}^{\ell}=\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=R_{k^{\ell}}^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell})).$
Substituting this definition into the hierarchical mixture, the overall function $f(\mathbf{x})$ becomes:
$f(\mathbf{x})=\sum_{k^{L}=1}^{K}R_{k^{L}}^{L}(f_{trans}^{L}(\mathbf{h}^{L}))g_{k^{L}}^{L}\Bigg{(}\dots\Bigg{(}\sum_{k^{1}=1}^{K}R_{k^{1}}^{1}(f_{trans}^{1}(\mathbf{h}^{1}))g_{k^{1}}^{1}(\mathbf{x})\Bigg{)}\Bigg{)}.$ (9)
This recursive formulation shows that at each layer, the submodules are weighted according to the input-dependent routing function, which adapts based on the current hidden state.
For a layer-wise perspective, the computation at each layer $\ell$ can be explicitly written as:
$\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}R_{k^{\ell}}^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))g^{\ell}_{k^{\ell}}(\mathbf{h}^{\ell}).$ (10)
Here, $R^{\ell}$ selects the most relevant submodules for a given input, allowing the network to adaptively compose its computation path at each layer in a sparse fashion.
3.2 Quantifying FLU with MixLoRA
In the previous section, we have established a parsimonious representation of the functional space. In this section, we choose a simple function to encode the functional level uncertainty and provide the intuition as follows. We first note that given a fixed MoE model architecture, the larger weight a mixture component has, the more certain we are about that component.
The model uncertainty of the MixLoRA architecture is quantified by considering perturbations $\Delta f(\mathbf{x})$ to the model $f(\mathbf{x})$. Following the discussion and notation in Sec. 3.1, we can show that these perturbations are instantiated in the space of the mixture weights $\alpha$ (as defined in Eq. 8 and 9):
Fact 3.1 (Model Perturbation Structure).
Under regularity assumptions on the residual connection architecture, perturbations $\Delta f(x)$ to the model $f(x)$ approximately decompose as:
$\Delta f(x)\approx\sum_{\ell=1}^{L}\sum_{k^{\ell}=1}^{K}\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell}).$
In high-dimensional settings, the basis functions are approximately orthogonal. Thus, the perturbation $\Delta f(x)$ is entirely represented by the set $\left\{\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\right\}_{k^{\ell}=1,\dots,K}^{\ell=1,\dots,L}$. Therefore, the functional-level uncertainty (FLU) can be generally modeled as a linear function over the mixture weights:
$\text{FLU}(x)=U\left(\left\{\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\right\}_{k^{\ell}=1,\dots,K}^{\ell=1,\dots,L}\right),$
where $U(\cdot)$ denotes a linear aggregation function. Details of derivation is presented in Appendix A.1.
In practice, as illustrated in Figure 2, the top-$k$ router at each layer produces a sparse probability vector $\alpha^{\ell}=(\alpha_{1}^{\ell},...,\alpha_{k}^{\ell})$, dynamically mixing the basis functions captured by the LoRA experts given the current hidden state. The values of the top-$K$ routing weights that contribute to the final output hidden state serve as a direct quantification of the model’s uncertainty at functional-level. We follow the routing mechanisms used in MoE layers (see Eq. (3) and (4)), employing top-$2$ gate routers for the mixture. At each layer $\ell$, we compute the raw router probabilities and retrain the largest two:
$\tilde{R}^{\ell}(\mathbf{h}^{\ell})=\text{Keep-Top-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}^{\ell})).$ (11)
Given an input prompt $x$ of length $s$, we aggregate the router weights over the selected experts and across all layers to estimate the FLU:
$\text{FLU}(x)=\frac{1}{L}\sum_{\ell=1}^{L}\sum_{i=1}^{2}\tilde{R}_{i}^{\ell}(\mathbf{h}^{\ell}).$ (12)
This formulation provides an efficient approach to quantify functional-level uncertainty in LoRA MoE architectures.
3.3 Calibration Loss